Explain why it is not necessary to state a subtraction principle: For any real numbers a, b, and c, a = b is equivalent to a – c = b – c.
Answer by stanbon(75887)

Explain why it is not necessary to state a subtraction principle: For any real numbers a, b, and c, a = b is equivalent to a – c = b – c.
-------------
One addition is defined for all Real Numbers,
subtraction is defined as addition of an
additive inverse.
a-b = a+(-b)
-------------------
Cheers,
Stan H.